OpenStudy (anonymous):

Find the derivative of the function f(x) = cos^2(x) + cot^2(x). 2cos(x) + 2cot(x) 2cos(x)sin(x) + 2cot^2(x) –2sin(x)cos(x) – 2cot(x)csc^2(x) 2cos(x)sin(x) – csc^2(x)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

what is the first derivative of \[{\left( {\cos x} \right)^2}\]

OpenStudy (anonymous):

-sin^2(x)

OpenStudy (anonymous):

No treat the 2 as a power rule

OpenStudy (anonymous):

You see Michele put the 2 in brackets which is the same thing as cos^2x just easier to see what's going on :), this required power + chain rule.

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

so it would be: (-sinx)^2

OpenStudy (michele_laino):

no I don't think, please keep in mind that \[{\left( {\cos x} \right)^2} = \cos x \cdot \cos x\] then apply the product rule

OpenStudy (anonymous):

so its -sinx times -sinx

OpenStudy (anonymous):

?:)

OpenStudy (michele_laino):

product rule is this: \[\frac{d}{{dx}}\left( {fg} \right) = \frac{{df}}{{dx}} \cdot g + f \cdot \frac{{dg}}{{dx}}\]

OpenStudy (anonymous):

\[f(x) = \cos^2x \implies~~f'(x) = 2(cosx)*cosx'\]

OpenStudy (anonymous):

Does that help? :P

OpenStudy (anonymous):

yes! (: i think the answer is c) –2sin(x)cos(x) – 2cot(x)csc2(x)

OpenStudy (michele_laino):

hint, if I apply the product rule, I can write: \[\frac{d}{{dx}}{\left( {\cos x} \right)^2} = \frac{{d\cos x}}{{dx}} \cdot \cos x + \cos x \cdot \frac{{d\cos x}}{{dx}} = ...?\]

OpenStudy (anonymous):

You got it math!

OpenStudy (anonymous):

thank u!!! :)

OpenStudy (michele_laino):

thank you!!

OpenStudy (anonymous):

I think you may have been confused by the notation by Michele, even though that is the "legit" way, but I use primes at the beginning since it's easier to see what's going on when people first start calculus at least :P